# Disentangling Task Interference within Neurons: Model Merging in Alignment with Neuronal Mechanisms

### v1

@@ Abstract
Fine-tuning pre-trained models on targeted datasets enhances task-specific performance but often comes at the expense of generalization. Model merging techniques, which integrate multiple fine-tuned models into a single multi-task model through task arithmetic at various levels: model, layer, or parameter, offer a promising solution. However, task interference remains a fundamental challenge, leading to performance degradation and suboptimal merged models. Existing approaches largely overlook the fundamental role of individual neurons and their connectivity, resulting in a lack of interpretability in both the merging process and the merged models. In this work, we present the first study on the impact of neuronal alignment in model merging. We decompose task-specific representations into two complementary neuronal subspaces that regulate neuron sensitivity and input adaptability. Leveraging this decomposition, we introduce NeuroMerging, a novel merging framework developed to mitigate task interference within neuronal subspaces, enabling training-free model fusion across diverse tasks. Through extensive experiments, we demonstrate that NeuroMerging achieves superior performance compared to existing methods on multi-task benchmarks across both vision and natural language domains. Our findings highlight the importance of aligning neuronal mechanisms in model merging, offering new insights into mitigating task interference and improving knowledge fusion.
Disentangling Task Interference within Neurons: Model Merging in Alignment with Neuronal Mechanisms
Zitao Fang1, Guodong Du2, Shuyang Yu1, Yifei Guo1, Yiwei Zhang1, Jing Li2, Ho-Kin Tang2*, Sim Kuan Goh1* 1Xiamen University Malaysia, 2Harbin Institute of Technology, Shenzhen, China ait2209071@xmu.edu.my, denghaojian@hit.edu.cn, simkuangoh@gmail.com
@@ 1 Introduction
Pre-trained models (PTMs), such as foundation models and large language models (LLMs) Vaswani et al. (2017); Achiam et al. (2023); Touvron et al. (2023), have revolutionized AI by learning rich representations from large-scale datasets. These models demonstrate general capabilities while enabling effective fine-tuning for task-specific adaptation Touvron et al. (2023). PTMs have driven significant advancements across core AI domains, including natural language processing (NLP), computer vision (CV), as well as applications in medicine, law, education Bommasani et al. (2021); Moor et al. (2023); Ray (2023). Building on this success, multi-task learning (MTL) has been a paradigm for integrating task-specific abilities into a model Fifty et al. (2021), allowing generalization across multiple specialized tasks. Nonetheless, MTL requires simultaneous training on all targeted datasets, which can be costly and pose privacy concerns. Model Merging Wortsman et al. (2022); Ilharco et al. (2023); Du et al. (2024) has recently emerged as an alternative paradigm to MTL for task adaptation, enabling the training-free integration of fine-tuned models, which are increasingly being shared publicly (e.g., on Hugging Face).
@@
Model Merging began with weight interpolation Wortsman et al. (2022) to combine the strengths of different models and has since evolved into techniques that balance competition and cooperation within shared representation and enable task editing in weight space Du et al. (2024); Ilharco et al. (2023); Ortiz-Jimenez et al. (2024). In NLP, methods such as merging task-specific language models have been explored to build and update foundation models with multi-task capabilities Raia (2021); Wan et al. (2024); Akiba et al. (2025); Wan et al. (2025). Similarly, in CV, approaches like merging Vision Transformers (ViTs) trained on different tasks or domains have been investigated to create unified models capable of handling diverse visual tasks Kim et al. (2021); Bao et al. (2022); Wang et al. (2024). In the multi-modal space, model merging has been applied to integrate models from different modalities, such as text and images, enhancing tasks like audio-visual question answering and image captioning Sung et al. (2023); Sundar et al. (2024); Dziadzio et al. (2024). These advancements underscore model merging as a promising avenue for future research.
Table 1: Model merging with different scales and granularity levels.
Method Scale Granularity Level
Fisher Merging[NeurIPS22] Fisher Matrix Parameter
RegMean[ICLR23] Inner Product Matrix Parameter
Task Arithmetic[ICLR23] Uniformed Task
Ties-Merging[NeurIPS23] Uniformed Parameter
DARE[ICML24] $1/(1-p)$ Parameter
LoraHub[COLM24] Evolver Searched Task
AdaMerging[ICLR24] Unsupervised Optimized Layer
PCB-Merging[NeurIPS24] Balancing Matrix Parameter
NeuroMerging (Ours) L1-Norm Neuron
Existing methods for model merging primarily operate at three granularities—model-level, layer-level, or parameter-level Ilharco et al. (2023); Yang et al. (2023); Du et al. (2024)—while overlooking the fundamental role of individual neurons and their connectivity Suhaimi et al. (2022); Stelzer et al. (2021), which underpins the learning process all neural networks from Perception Rosenblatt (1958) to LLMs Touvron et al. (2023). In Figure 1, we illustrate that modifying model weights along two complementary neuronal subspaces—by removing one and retaining the other—leads to distinct impacts on task performance. Notably, one subspace preserves most of the task-specific capabilities. This observation motivates us to explore model merging at the neuronal level, which could have important implications for mitigating task interference and could yield more robust merged models.
In this work, we present the first study to examine task interference at the neuronal level. Specifically, we investigate the role of neuronal alignment in model merging, illustrated in Figure 2. We begin by decomposing task-specific representations into two complementary neuronal subspaces that regulate neuron sensitivity and input adaptability. Leveraging the insights from the decomposition, we introduce NeuroMerging, a novel merging framework designed to mitigate task interference within neuronal subspaces, enabling training-free model fusion across diverse tasks. To evaluate our approach, we conduct experiments on multi-task benchmarks across both vision and natural language domains, considering various settings, including in-domain and out-of-domain generalization. Empirically, our method outperforms existing approaches. The main contributions of our paper are as follows:
  • We present the first exploration into the impact of neuronal alignment in the merging process, introducing a decomposition of task-specific representations into two complementary neuronal subspaces.
  • Based on the insights from the neuronal subspaces, we propose NeuroMerging, a new framework developed to reduce task interference in alignment with neuronal mechanism.
  • We show that NeuroMerging achieved superior performance compared to existing approaches on multi-task benchmarks in both vision and natural language processing.
2 Related Work
@@
Multi-task learning (MTL) Fifty et al. (2021) leverages transferable knowledge to handle multiple related tasks simultaneously. Existing MTL approaches primarily rely on architectural design or optimization strategies. Architectural-based methods, such as Mixture of Experts (MoE) Shazeer et al. (2017), introduce specialized subnetworks that dynamically route inputs to task-specific experts, effectively reducing interference. However, these methods require modifying the pretrained model structure, increasing computational complexity, and limiting scalability Liu et al. (2019); Shen et al. (2024). Optimization-based approaches, on the other hand, focus on balancing task gradients or loss functions to mitigate task conflicts during training Bai et al. (2023); Kendall et al. (2018). While these methods improve convergence, they still depend on task-specific training data, which may be impractical in real-world applications due to privacy concerns or data scarcity Liang et al. (2020). In contrast, model merging offers an alternative paradigm by integrating knowledge from multiple fine-tuned models into a single unified model without requiring additional training data or architectural modifications Wortsman et al. (2022); Ilharco et al. (2023). Notwithstanding the promising findings, a key challenge in model merging is task conflict Yadav et al. (2024); Du et al. (2024), where different tasks compete for model capacity, potentially leading to suboptimal performance.
@@
To resolve task conflicts, existing model merging methods can be categorized into three levels based on their granularity. Model-level merging combines entire model weights, typically through averaging or weighted aggregation, but often results in performance degradation due to the loss of task-specific knowledge, as seen in methods like Task Arithmetic Ilharco et al. (2023) and LoRAHub Huang et al. (2023). Layer-level merging selectively integrates layers from different models under the assumption of shared representations; for instance, AdaMerging Yang et al. (2023) adapts layer selection to better preserve task-specific information. Parameter-level merging directly manipulates individual parameters to blend knowledge from multiple models, enhancing adaptability and robustness. Techniques such as Fisher Merging Matena and Raffel (2022), RegMean Jin et al. (2023), TIES-Merging Yadav et al. (2024), DARE Le et al. (2024), and PCB-Merging Du et al. (2024) exemplify this approach. However, existing methods largely overlook the fundamental role of neuron and neuron-level interactions in task specialization. In this work, we aim to bridge this gap.
@@
Current neural network models, from the Perceptron invented by Rosenblatt (1958) to recent massive LLMs Touvron et al. (2023), have grown significantly in scale and complexity. Nevertheless, the core principle remains unchanged: individual neurons and their connectivity still underpin the learning process Stelzer et al. (2021); Suhaimi et al. (2022). During the pre-training and fine-tuning process, neurons are not merely passive components but active elements of the network, each contributing to learning and inference Jiang et al. (2024); Islam et al. (2023). In this work, we conduct the first in-depth study on neuronal alignment in model merging, aiming to uncover its role in preserving task-specific knowledge and its potential to mitigate task conflict.
@@ 3 Methodology
In this section, we first formalize the concept of neuronal task vector for model merging and then decompose neuronal task vectors into two complementary neuronal subspaces. Subsequently, we introduce our framework, NeuroMerging, showed in Algorithm 1, which performs merging in the neuronal subspaces.
@@ 3.1 Preliminaries
In this work, we consider the model merging of a set of $T$ task specific models, $(\theta_{1},\dots,\theta_{t},\dots,\theta_{T})$, fine-tuned from a pretrained model $\theta_{0}$. With the task vector notation, each task is defined as $\tau_{t}=\theta_{t}-\theta_{0}$. The merged model is $\bar{\theta}=\theta_{0}+\phi(\tau_{1},...,\tau_{t},...,\tau_{T})$, where $\phi(\cdot)$ represents the transformation applied to each task vector $\tau_{t}$ and followed by merging. As entries of the task vector with larger magnitudes are more relevant to task-specific adaptation, only the top $r\%$ of $\mathbf{\tau}_{t}$ with the largest magnitudes are kept, while the others are set to zero Yadav et al. (2024). The masked task vector is defined as $\tau^{\text{masked}}_{t}=m_{t}\circ\tau_{t}$, where $m_{t}$ is the mask that keeps the top $r\%$ of the elements of each task vector. In the following section, we use $\tau_{t}$ to represent the masked task vector for readability.

### v2

# To See a World in a Spark of Neuron: Disentangling Multi-task Interference for Training-free Model Merging

@@ Abstract
Fine-tuning pre-trained models on targeted datasets enhances task-specific performance but often comes at the expense of generalization. Model merging techniques, which integrate multiple fine-tuned models into a single multi-task model through task arithmetic, offer a promising solution. However, task interference remains a fundamental challenge, leading to performance degradation and suboptimal merged models. Existing approaches largely overlook the fundamental roles of neurons, their connectivity, and activation, resulting in a merging process and a merged model that does not consider how neurons relay and process information. In this work, we present the first study that relies on neuronal mechanisms for model merging. We decompose task-specific representations into two complementary neuronal subspaces that regulate neuron sensitivity and input adaptability. Leveraging this decomposition, we introduce NeuroMerging, a novel merging framework developed to mitigate task interference within neuronal subspaces, enabling training-free model fusion across diverse tasks. Through extensive experiments, we demonstrate that NeuroMerging achieves superior performance compared to existing methods on multi-task benchmarks across both natural language and vision domains. Our findings highlight the importance of aligning neuronal mechanisms in model merging, offering new insights into mitigating task interference and improving knowledge fusion. Code will be released upon acceptance.
To See a World in a Spark of Neuron: Disentangling Multi-task Interference for Training-free Model Merging
Zitao Fang1, Guodong Du2, Shuyang Yu1, Yifei Guo1, Yiwei Zhang1, Yiyao Cao1, Jing Li2, Ho-Kin Tang2*, Sim Kuan Goh1* 1Xiamen University Malaysia, 2Harbin Institute of Technology, Shenzhen, China ait2209071@xmu.edu.my, denghaojian@hit.edu.cn, simkuangoh@gmail.com
@@ 1 Introduction
"To see a world in a grain of sand…"
–William Blake
Figure 1: Comparative analysis of our method with baselines across models and domains.
Aspect Task Vector Neuronal Task Vector
Vector Space Resides in the weight space Resides in two complementary neuronal subspaces of neurons
Interpretation Represents weight interpolation between pre-trained and fine-tuned models Aligns with neuronal mechanisms, with input sensitivity and task adaptability
Activation $y=\phi\left(\left(\mathbf{w}_{0}^{k}+\tau_{t}^{k}\right)\cdot\bar{\mathbf{x}}\right)$ $y=\phi\left(\mathbf{w}_{0}^{k}\cdot\bar{\mathbf{x}}_{\parallel}+\tau_{\parallel,t}^{k}\cdot\bar{\mathbf{x}}_{\parallel}+\tau_{\perp,t}^{k}\cdot\bar{\mathbf{x}}_{\perp}\right)$
Table 1: Differences between task vector and neuronal task vector. Please refer to Section 3 for details.
To see a world in the spark of a neuron, the essence of deep learning lies in the complex dynamics of interconnected neurons, which has been empowering recent pre-trained models (PTMs), such as foundation models and large language models (LLMs) Vaswani et al. (2017); Achiam et al. (2023); Touvron et al. (2023), to learn rich representations from large-scale datasets. These models demonstrate general capabilities while enabling effective fine-tuning for task-specific adaptation Touvron et al. (2023). PTMs have driven significant advancements across core AI domains, including natural language processing (NLP), computer vision (CV), as well as applications in medicine, law, education Bommasani et al. (2021); Moor et al. (2023); Ray (2023). Building on this success, multi-task learning (MTL) has been a paradigm for integrating task-specific abilities into a model Fifty et al. (2021), allowing generalization across multiple specialized tasks. Nonetheless, MTL requires simultaneous training on all targeted datasets, which can be costly and pose privacy concerns. Model Merging Wortsman et al. (2022); Ilharco et al. (2023); Du et al. (2024) has recently emerged as an alternative paradigm to MTL for task adaptation, enabling the training-free integration of fine-tuned models, which are increasingly being shared publicly (e.g., on Hugging Face).
Model Merging began with weight interpolation Wortsman et al. (2022) to combine the strengths of different models in weight space by balancing competition and cooperation within shared representation Du et al. (2024); Ilharco et al. (2023); Ortiz-Jimenez et al. (2023). In NLP, methods such as merging task-specific language models have been explored to build and update foundation models with multi-task capabilities Raia (2021); Wan et al. (2024a); Akiba et al. (2025); Wan et al. (2024b). Similarly, in CV, approaches like merging Vision Transformers (ViTs) trained on different tasks or domains have been investigated to create unified models capable of handling diverse visual tasks Kim et al. (2021); Bao et al. (2022); Wang et al. (2024). In the multi-modal space, model merging has been applied to integrate models from different modalities, such as text and images, enhancing tasks like audio-visual question answering and image captioning Sung et al. (2023); Sundar et al. (2024); Dziadzio et al. (2025). These advancements underscore model merging as a promising avenue for future research.
Existing methods for model merging primarily operate at three granularities: model-level, layer-level, or parameter-level Ilharco et al. (2023); Yang et al. (2024b); Du et al. (2024), while overlooking the fundamental roles of neurons, their activation and connectivity Suhaimi et al. (2022); Stelzer et al. (2021), which underpins the learning process all neural networks from Perception Rosenblatt (1958) to LLMs Touvron et al. (2023). Appendix Table 6 provided a summary of existing methods across different granularities. In Figure 2 and Appendix Table 6.2, we analyze on how changes in model weights along two complementary neuronal subspaces, leads to distinct impacts on task performance, by removing one and retaining the other. Notably, one subspace preserves most of the task-specific capabilities. This observation motivates us to explore model merging at the neuronal level, which could have important implications for mitigating task interference and could yield more robust merged models.
@@
In this work, we present the first study to examine model merging at the neuronal level, illustrated in Figure 3. Specifically, we investigate the roles of neuronal mechanisms (i.e., neuron connectivity and activation) in model merging, both mathematically and empirically. We begin by decomposing task-specific representations from fine-tuned models into two complementary neuronal subspaces. These subspaces are mathematically characterized to show that they regulate neuron sensitivity and input adaptability. The key differences of our neuronal subspaces with existing task vectors summarized in Table 1. Leveraging the insights from the decomposition, we introduce NeuroMerging, a novel merging framework designed to mitigate task interference within neuronal subspaces, enabling training-free model fusion across diverse tasks. To evaluate our approach, we conduct experiments on multi-task benchmarks across both natural language and vision domains, considering various settings, including in-domain and out-of-domain generalization. Empirically, our method outperforms existing approaches as summarized in Figure 1. The main contributions of our paper are as follows:
  • We presented the first exploration into the roles of neuronal mechanisms in the merging process, introducing a decomposition of task-specific representations into two complementary neuronal subspaces.
  • Based on the insights from the neuronal subspaces, we proposed NeuroMerging, a new framework developed for model merging accounting input sensitivity and task specificity.
  • We showed that NeuroMerging achieved superior performance compared to existing approaches on multi-task benchmarks in both natural language processing and vision domains.
2 Related Work
@@
Multi-task learning (MTL) Fifty et al. (2021) leverages transferable knowledge to handle multiple related tasks simultaneously. Existing MTL approaches primarily rely on architectural design or optimization strategies. Architectural-based methods, such as Mixture of Experts (MoE) Shazeer et al. (2017), introduce specialized subnetworks that dynamically route inputs to task-specific experts, effectively reducing interference. However, these methods require modifying the pretrained model structure, increasing computational complexity, and limiting scalability Liu et al. (2019); Shen et al. (2024). Optimization-based approaches, on the other hand, focus on balancing task gradients or loss functions to mitigate task conflicts during training Bai et al. (2023); Cipolla et al. (2018). While these methods improve convergence, they still depend on task-specific training data, which may be impractical in real-world applications due to privacy concerns or data scarcity Liang et al. (2022). In contrast, model merging offers an alternative paradigm by integrating knowledge from multiple fine-tuned models into a single unified model without requiring additional training data or architectural modifications Wortsman et al. (2022); Ilharco et al. (2023). Notwithstanding the promising findings, a key challenge in model merging is task conflict Yadav et al. (2023); Du et al. (2024), where different tasks compete for model capacity, potentially leading to suboptimal performance.
@@
To resolve task conflicts, existing model merging methods can be categorized into three levels based on their granularity. Model-level merging combines entire model weights, typically through averaging or weighted aggregation, but often results in performance degradation due to the loss of task-specific knowledge, as seen in methods like Task Arithmetic Ilharco et al. (2023) and LoRAHub Huang et al. (2024a). Layer-level merging selectively integrates layers from different models under the assumption of shared representations; for instance, AdaMerging Yang et al. (2024b) adapts layer selection to better preserve task-specific information. Parameter-level merging directly manipulates individual parameters to blend knowledge from multiple models, enhancing adaptability and robustness. Techniques such as Fisher Merging Matena and Raffel (2022), RegMean Jin et al. (2023), TIES-Merging Yadav et al. (2023), DARE Yu et al. (2024a), and PCB-Merging Du et al. (2024) exemplify this approach. However, existing methods largely overlook the fundamental role of neuron and neuron-level interactions in task specialization. In this work, we aim to bridge this gap.
@@
Current neural network models, from the Perceptron invented by Rosenblatt (1958) to recent massive LLMs Touvron et al. (2023), have grown significantly in scale and complexity. Nevertheless, the core principle remains unchanged: neurons and their connectivity still underpin the learning process Stelzer et al. (2021); Suhaimi et al. (2022). During the pre-training and fine-tuning process, neurons are not merely passive components but active elements of the network, each contributing to learning and inference Jiang et al. (2024); Islam et al. (2023). In this work, we attempt to conduct the first in-depth study on neuronal mechanisms in model merging.
@@ 3 Methodology
In this section, we first formalize the concept of neuronal task vector for model merging and then decompose neuronal task vectors into two complementary neuronal subspaces. Subsequently, we introduce our framework, NeuroMerging, showed in Algorithm 1 in Appendix A, which performs merging in the neuronal subspaces.
@@ 3.1 Preliminaries
Model merging considers the fusion of a set of $T$ task specific models, $(\theta_{1},\dots,\theta_{t},\dots,\theta_{T})$, fine-tuned from a pretrained model $\theta_{0}$. With the task vector notation, each task is defined as $\tau_{t}=\theta_{t}-\theta_{0}$. The merged model is $\bar{\theta}=\theta_{0}+\xi(\tau_{1},...,\tau_{t},...,\tau_{T})$, where $\xi(\cdot)$ represents the transformation applied to each task vector $\tau_{t}$ and followed by merging.
Zooming into the neuronal level of $\tau_{t}$ from any neural network models (from perceptron to LLMs), the activation of the neuron can be computed with an input $\bar{x}\in\mathbb{R}^{n}$:
$y=\phi(\mathbf{w}^{k}_{t}\cdot\bar{x})$ (1)
where $\mathbf{w}^{k}_{t}$ represents the synaptic connectivity to the $k^{th}$ neuron after fine-tuning from the pre-trained weights $\mathbf{w}^{k}_{0}$. $\phi(\cdot)$ denotes non-linear activation function. Noting that although the input $\bar{x}$ is involved in the definition and derivation, it is not required during the merging process.
3.2 Neuronal Task Vector
Neuronal task vector is defined to be the difference between task-specific fine-tuning of a neuron with the pre-trained model: $\mathbf{\tau}^{k}_{t}=\mathbf{w}^{k}_{t}-\mathbf{w}^{k}_{0}$. Based on the defined neuronal task vector, we can rewrite the activation of the $k^{th}$ neuron as:
$y=\phi((\mathbf{w}^{k}_{0}+\tau^{k}_{t})\cdot\bar{x})$ (2)
3.3 Neuronal Subspace Decomposition
To examine how task-specific fine-tuning impact neurons, we decompose the neuronal task vectors into two complementary neuronal subspaces, visualized in Figure 3. Mathematically, this decomposition is formulated as:
$\tau^{k}_{t}=\tau^{k}_{\parallel,t}+\tau^{k}_{\perp,t},$ (3)
where $\tau^{k}_{\parallel,t}=\mathbf{P}\tau^{k}_{t}$ projects the neuronal task vectors onto the pre-trained model’s weight space, Parallel Subspace ($\mathcal{P}$). Here, $\mathbf{P}$ is the projection matrix onto the span of $\mathbf{W}^{k}_{0}$. $\tau^{k}_{\perp,t}=(\mathbf{I}-\mathbf{P})\tau^{k}_{t}$ captures the complementary orthogonal modifications, in the Orthogonal Subspace ($\mathcal{O}$). Noting that $\bar{x}$ is not required during decomposition.
Based on Neuronal Subspace Decomposition, we decompose both $\tau^{k}_{t}$ and $\bar{x}$ into two orthogonal subspaces ($\mathbf{P}$ and $\mathcal{O}$) of $\mathbf{w}^{k}_{0}$ and obtain:
$y=\phi((\mathbf{w}^{k}_{0}+\tau^{k}_{\parallel,t}+\tau^{k}_{\perp,t})\cdot(\bar{x}_{\parallel}+\bar{x}_{\perp}))$ (4)
Due to the orthogonality between vectors in the parallel subspace and those in the orthogonal subspace, we can simplify the expression as:
$y=\phi(\mathbf{w}^{k}_{0}\cdot\bar{x}_{\parallel}+\tau^{k}_{\parallel,t}\cdot\bar{x}_{\parallel}+\tau^{k}_{\perp,t}\cdot\bar{x}_{\perp})$ (5)
Since $\mathbf{w}^{k}_{0}$ and $\tau^{k}_{\parallel,t}$ are parallel, we can further express as:
$y=\phi(s(\tau^{k}_{\parallel,t})\cdot\bar{x}_{\parallel}+\tau^{k}_{\perp,t}\cdot\bar{x}_{\perp})$ (6)
Thus, we obtain a mechanistic view of the fine-tuned neuron: it adjusts its sensitivity $s$ to input in the parallel subspace ($\mathcal{P}$) while adapting itself to capture $\bar{x}_{\perp}$ to handle task $t$ in the orthogonal subspace ($\mathcal{O}$).
The role of each complementary subspace:
3.4 NeuroMerging
Subsequently, we reconstruct the task vector $\overline{\tau}$ with the all the merged neuronal task vectors $\overline{\tau}_{k}$. Finally, we obtained the final merged model $\overline{\theta}=\theta_{0}+\overline{\tau}_{rescaled}$. Additional discussion is provided in Appendix B.
4 Experimental Setup
Baseline Methods. Our baselines comprise two main categories: (1) non-model merging approaches, which include individually fine-tuned models and a multitask model trained jointly on the combined dataset serving as our theoretical upper bound, and (2) various advanced model merging techniques, including Simple Averaging (Wortsman et al., 2022), Fisher Merging (Matena and Raffel, 2022), RegMean (Jin et al., 2023), Task Arithmetic (Ilharco et al., 2023), TIES-Merging (Yadav et al., 2023), AdaMerging (Yang et al., 2024b), PCB-Merging (Du et al., 2024), and our proposed NeuroMerging method. Notably, we use task-wise AdaMerging for fair comparison with existing baselines and our method. We report average accuracy across all tasks’ test sets as our primary evaluation metric.
Task($\rightarrow$) Test Set Performance
Zeroshot - 53.1 58.2 54.2 54.1 54.3 70.9 49.2 63.9
Finetuned - 88.0 94.4 97.1 85.3 91.0 95.7 71.6 80.6
Multitask - 88.1 94.2 98.5 89.3 92.0 95.4 73.5 73.6
Averaging[ICML22] 51.6 59.2 26.3 69.6 53.8 67.3 49.1 36.1
Task Arithmetic[ICLR23] 59.7 60.9 31.7 57.8 73.0 73.5 55.7 65.3
TIES-Merging[NeurIPS23] 76.7 80.8 92.4 77.7 81.9 78.4 61.9 63.9
PCB-Merging[NeurIPS24] 76.9 82.9 93.2 79.0 84.4 75.6 63.5 59.7
NeuroMerging (Ours) 77.6 81.1 94.3 81.6 84.7 81.2 56.7 83.9
Fisher Merging[NeurIPS22] 68.7 68.4 83.0 65.5 62.4 94.1 58.2 49.2
RegMean[ICLR23] 79.8 83.9 97.2 73.2 82.6 94.1 63.2 64.4
Task Arithmetic[ICLR23] 74.6 72.7 91.3 76.4 85.6 74.4 61.0 61.1
TIES-Merging[NeurIPS23] 79.5 82.6 94.9 72.8 87.4 85.2 66.6 66.7
PCB-Merging[NeurIPS24] 81.0 87.0 95.2 76.4 88.1 88.4 64.3 68.1
NeuroMerging (Ours) 82.2 86.4 94.3 75.9 87.9 91.2 65.9 73.6
Task vector-based approaches such as Task Arithmetic, Ties-Merging, and PCB-Merging, along with our proposed NeuroMerging, are substantially more lightweight and efficient. These approaches are training-free and do not rely on a validation set, making them highly practical for real-world applications. To further evaluate this advantage, we conducted additional experiments comparing task vector-based methods in scenarios where validation sets were unavailable.
5 Results
Figure 4: In-domain v.s. Out-domain performance.
5.1 Merging NLP Models
Following the experimental settings from Yadav et al. (2023), we use the T5-base and T5-large models (Raffel et al., 2020), which are encoder-decoder transformers (Vaswani et al., 2017) pretrained via masked language modeling on a large text corpus, and fine-tune them independently on seven tasks: Khot et al.’s (2019) QASC, Yang et al.’s (2015) WikiQA, and Tafjord et al.’s (2019) QuaRTz for Question Answering; Zhang et al.’s (2019) PAWS for Paraphrase Identification; Sharma et al.’s (2018) Story Cloze for Sentence Completion; and Sakaguchi et al.’s (2021) Winogrande together with Levesque et al.’s (2012) WSC for Coreference Resolution. Table 2 and 12 demonstrated that our approach has superior performance over state-of-the-art methods, achieving improvements of $0.6\%$ and $1.2\%$ for T5-base and T5-large, respectively. Moreover, NeuroMerging without validation showed even more substantial gains, surpassing previous methods by $1.4\%$ for T5-base and $0.7\%$ for T5-large. For comprehensive results across all tasks and model variants, see Appendix C.3.
5.2 Out-of-Domain Generalization
Building upon the experimental setup of Yadav et al. (2023), we also investigated how merged models between tasks enhance generalization in different domains. Following the approach used in prior NLP models, we merged models on seven in-domain datasets and evaluated their performance on six held-out datasets from the T0 mixture (Sanh et al., 2022) to assess cross-task generalization. These datasets encompass diverse tasks, covering three Question Answering datasets: Huang et al.’s (2019) Cosmos QA, Sap et al.’s (2019) Social IQA, and Rogers et al.’s (2020) QuAIL; one Word Sense Disambiguation dataset: Pilehvar and Camacho-Collados’s (2019) WiC; and two Sentence Completion datasets: Gordon et al.’s (2012) COPA and Zellers et al.’s (2019) H-SWAG. As shown in Figure 4, NeuroMerging outperformed the strongest baseline by $0.7\%$ and $0.5\%$ for T5-base and T5-large models, respectively, showcasing enhanced out-of-domain generalization capabilities. For more comprehensive results, please refer to Appendix C.3 Table 15 and 16.
5.3 Merging LLMs
Model CMMLU GSM8K Human-Eval Average
Chinese 38.6 2.3 13.4 18.1
Math 31.2 65.6 0.0 32.3
Code 33.3 0.0 17.1 16.8
Averaging[ICML22] 35.6 48.5 6.7 30.3
Task Arithmetic[ICLR23] 35.4 46.1 9.8 30.4
TIES-Merging[NeurIPS23] 36.5 53.4 12.8 34.3
PCB-Merging[NeurIPS24] 36.4 52.3 16.5 35.1
PCB-Merging+ES [NeurIPS24] 36.4 53.1 16.5 35.3
NeuroMerging (Ours) 36.1 57.2 14.6 36.0
Table 3: Performance comparison on LLaMA2.
We follow the experimental setup of Du et al. (2024) and extend our NeuroMerging to larger LLMs. Specifically, we merged three domain-specialized Llama-2-7b models (Touvron et al., 2023), and each is fine-tuned for distinct capabilities: Chinese language understanding, mathematical reasoning (Yu et al., 2024b), and code generation (Roziere et al., 2023). To rigorously evaluate the performance of each specialized model, we employed established benchmarks tailored to their respective domains: Li et al.’s (2024) CMMLU for assessing Chinese language proficiency, Cobbe et al.’s (2021) GSM8K for mathematical capabilities, and Chen et al.’s (2021) HumanEval for code generation competency. As demonstrated in Table 3, our NeuroMerging approach exhibited substantial performance improvements, surpassing the strongest baseline by $0.9\%$. Notably, our method even outperformed the PCB-Merging utilizing Evolution Strategies (ES) optimization algorithm by $0.7\%$, underscoring the effectiveness of our proposed methodology.
5.4 Merging Vision Models
Method Validation ViT-B/32 Avg. ViT-L/14 Avg.
Individual - 90.5 94.2
Multi-task - 88.9 93.5
Averaging[ICML22] 65.8 79.6
Task Arithmetic[ICLR23] 60.4 83.3
TIES-Merging[NeurIPS23] 72.4 86.0
PCB-Merging[NeurIPS24] 75.9 86.9
NeuroMerging (Ours) 76.4 87.9
Fisher Merging[NeurIPS22] 68.3 82.2
RegMean[ICLR23] 71.8 83.7
Task Arithmetic[ICLR23] 70.1 84.5
TIES-Merging[NeurIPS23] 73.6 86.0
AdaMerging[ICLR2024] 71.1 84.9
AdaMerging++[ICLR2024] 73.7 87.3
PCB-Merging[NeurIPS24] 76.3 87.5
NeuroMerging (Ours) 76.5 88.3
Table 4: Performance comparison on ViT.
6 Additional Results and Analysis
6.1 Merging without Validation Sets
When validation data is unavailable, we examine the parameters $\lambda_{1}$ and $\lambda_{2}$ selected according to Section 3.4, where $\lambda_{1}$ is set to zero, as the corresponding subspace is observed to have little impact in Section 6.3. The value of $\lambda_{2}$ is computed based on the L1-norm of masked and unmasked task vectors. Figure 8 and Table 2 and Appendix Tables 2,  13, and 14 present the evaluation of NeuroMerging on NLP and CV tasks across various model sizes, comparing it with existing methods. NeuroMerging achieves the highest average accuracy across all tasks. This demonstrates that our proposed method, with a simple rescaling, outperforms existing methods on average. Specifically, it achieved a 1.4% and 0.7% improvement over the strongest baseline for T5-Base and T5-Large, respectively. For vision models, it outperforms the strongest baseline by 0.5% and 1.0% for ViT-B/32 and ViT-L/14, respectively.
6.2 Role of Neuronal Subspaces
T5-large In-domain Out-domain Total Average
Fine-tuned 88.0 53.8 58.7
Keep Orthogonal 88.0 53.9 58.8
Keep Parallel 52.7 51.8 51.9
Table 5: In/Out-domain Performance of T5-large.
6.3 Robustness of Hyperparameters
Figure 5: Impacts of $\lambda_{1}$ and $\lambda_{2}$ on T5-Large.
7 Conclusion
In this paper, we revisited model merging from the core principle neuron connectivity and activation that underpin the learning process of recent deep neural networks and LLMs. Specifically, we presented the first exploration into the roles of neuronal mechanisms in the model merging process, by decomposing of task-specific representations into two complementary neuronal subspaces for characterization. It is mathematically shown that neuronal subspaces regulate input sensitivity and task adaptability. Based on these insights, we proposed NeuroMerging, a novel framework designed to reduce task interference within neurons. Our empirical evaluations demonstrate that NeuroMerging achieves superior performance compared to existing approaches on multi-task benchmarks across both vision and natural language processing tasks.
Future work could extend NeuroMerging to larger models or multimodal architectures with more tasks. While our study introduces a neuronal mechanistic perspective for in model merging, it focuses only on neuron connectivity and activation, without considering higher-order neuronal dynamics (e.g., network level connectivity and dynamics). Further research is needed to investigate these factors and deepen our understanding of neuronal mechanisms in model merging, or even in multi-task learning.
Limitations
While our work provides the first neuronal mechanistic perspective on multi-task interference when merging large models, (1) it remains a partial view of neuronal mechanisms as it does not yet explore higher order neuronal dynamics during merging and inference, which require further investigation. Moreover, this work shares similar limitations with current SOTA model merging methods, including (2) the effectiveness of task arithmetic in model merging relies on selecting fine-tuned checkpoints that are beneficial for specific domains, ensuring they originate from the same pretrained model, and addressing hyperparameter sensitivity; and (3) More effort is needed to develop a mathematical understanding of why and when task arithmetic in model merging works well, despite its simplicity and efficiency.
Ethical Considerations
Our research is based on publicly available datasets and models, all of which conform to their respective licenses and ethical guidelines. While the proposed NeuroMerging approach itself introduces no immediate ethical risks, caution should be exercised when generalizing results beyond the evaluated domains. Specifically, applying this technique to privacy-sensitive or high-stakes scenarios may require additional validation, as performance and reliability in untested contexts remain uncertain. We thus recommend thorough evaluation and responsible deployment practices in such applications.
Appendix A Additional Details